# Cryogenic platform to investigate strong microwave cavity-spin coupling in correlated magnetic materials

### 07 Dec 2023 v1

@@ 3.4 Interpretation
Overall, the primary difference between the two sample spectra studied down to K is that CuSOHO is seen to strongly hybridize with the cavity mode, increasing the degree of hybridization with a decreasing temperature, while MgCrO shows no hybridization with the resonator environment despite experimental considerations such as filling factor, resonance frequency, and mode-volume confinement being the same.
@@
The coupling strength between paramagnetic spins and the microwave cavity is provided as [3]:
@@
where is the number of magnetic ions in the sample and is the cavity mode volume. For CuSOHO at K, the coupling strength is estimated to be on the order of MHz, but a more precise estimate would require modeling the spin-cavity Hamiltonian. The lack of hybridization for MgCrO is speculated to arise from its more strongly correlated paramagnetic spins that produce a broad ESR response even at room temperature, however, the mere observation of sample-cavity coupling differences between these two samples demonstrates that our setup probes both the hybridization and magnetic response across a wide range of temperatures in a relatively simple experimental architecture.
4 Conclusion
@@
This manuscript reports the design and implementation of loop-gap resonators (LGRs) for electron spin resonance (ESR) on quantum materials and hybrid quantum system studies of hybrid microwave polariton to quantum spin hybridization. The experimental setup is low-cost and implemented in cryogenic equipment commonly found in laboratories worldwide.

### 31 Jan 2024 v2

@@ III.4 Interpretation
Overall, the primary difference between the two sample spectra studied down to K is that CuSOHO is seen to strongly hybridize with the cavity mode, increasing the degree of hybridization with a decreasing temperature. At the same time, MgCrO shows no hybridization with the resonator environment despite experimental considerations such as filling factor, resonance frequency, and mode-volume confinement being the same.
@@
An expression showing the coupling strength between antiferromagnetic spins and the microwave cavity is provided by [3]:
where is the number of magnetic ions in the sample, is the cavity mode volume, and is the filling factor. This expression illustrates the dependence of coupling strength on the ratio of cavity resonance frequency () to mode volume characteristic of microwave resonators and explains the necessity of small-volume LGRs. To approximate the coupling strength for CuSOHO at K, the 1st derivative of with respect to the field () is taken and then plotted in a fashion similar to [28]. The results are shown in Fig. 5(a). Taking the derivative of in this way helps to remove the constant signal created by the coax cable and allows a more precise determination of the avoided level crossing. With a large part of the coax signal removed in the derivative, numerically integrating the derivative trace at Oe gives a decent approximation of the vacuum Rabi splitting, as shown in Fig. 5(b). The gap between peaks is given as MHz. The asymmetry in the peaks is likely due to unwanted features from the coax signal that still influence the reconstructed data. Based on this figure, the coupling strength is estimated to be to MHz, with a more precise estimate requiring experimental improvements to eliminate the cable resonances or modeling the spin-cavity Hamiltonian.
@@
The lack of hybridization for MgCrO is speculated to arise from its more strongly correlated paramagnetic spins that produce a broad ESR response even at room temperature. However, observing sample-cavity coupling differences between these two samples demonstrates that our setup probes both the hybridization and magnetic response across a wide range of temperatures in a relatively simple experimental architecture.
IV Conclusion
@@
This manuscript reports the design and implementation of loop-gap resonators (LGRs) for electron spin resonance (ESR) on quantum materials and hybrid quantum system studies using microwave polariton modes. The experimental setup is low-cost and implemented in cryogenic equipment commonly found in laboratories worldwide.
@@
Acknowledgements.
The work at Georgia Tech was supported by the National Science Foundation under award NSF-DMR-1750186 (A.K.J., M.M.). This work was performed, in part, at the Center for Integrated Nanotechnologies, an Office of Science User Facility operated for the U.S. Department of Energy (DOE) Office of Science. Sandia National Laboratories is a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia, LLC, a wholly owned subsidiary of Honeywell International, Inc., for the U.S. DOE’s National Nuclear Security Administration under contract DE-NA-0003525. The views expressed in the article do not necessarily represent the views of the U.S. DOE or the United States Government. We are grateful to Oleg Starykh for nudging us into the realm of ESR in quantum magnets and Henrik Rønnow for pointing out invaluable references to us.References
- [1] van der Est A 2016 Continuous-wave EPR vol 5 pp 1411–1422 ISBN 9780470058213
- [2] Ball J R, Yamashiro Y, Sumiya H, Onoda S, Ohshima T, Isoya J, Konstantinov D and Kubo Y 2018 Applied Physics Letters 112 204102 ISSN 0003-6951 publisher: American Institute of Physics URL https://aip.scitation.org/doi/10.1063/1.5025744
- [3] Everts J R, King G G G, Lambert N J, Kocsis S, Rogge S and Longdell J J 2020 Physical Review B 101 214414
- [4] Libersky M, McKenzie R, Silevitch D, Stamp P and Rosenbaum T 2021
- [5] Yang Y 2023 Investigation of Electro-Nuclear Spin States in LiHoF4 Using Cavity-Magnon-Polariton Technique Ph.D. thesis EPFL Lausanne
- [6] Oshikawa M and Affleck I 1999 Phys. Rev. Lett. 82(25) 5136–5139 URL https://link.aps.org/doi/10.1103/PhysRevLett.82.5136
- [7] Povarov K Y, Smirnov A I, Starykh O A, Petrov S V and Shapiro A Y 2011 Physical Review Letters 107 037204
- [8] Luo Z X, Lake E, Mei J W and Starykh O A 2018 Physical Review Letters 120 037204
- [9] Sichelschmidt J, Schlender P, Schmidt B, Baenitz M and Doert T 2019 Journal of Physics: Condensed Matter 31 205601 URL https://dx.doi.org/10.1088/1361-648X/ab071d
- [10] Joshi G, Kubasek J, Nikolov I, Sheehan B, Costa T A, Allão Cassaro R A and Friedman J R 2020 Review of Scientific Instruments 91 023104 ISSN 0034-6748 publisher: American Institute of Physics URL https://aip.scitation.org/doi/10.1063/1.5133074
- [11] Wood R L, Froncisz W and Hyde J S 1984 Journal of Magnetic Resonance (1969) 58 243–253 ISSN 00222364 URL https://linkinghub.elsevier.com/retrieve/pii/0022236484902142
- [12] Bissey J C, Berger R and Servant Y 1995 Solid State Communications 93 243–247 ISSN 0038-1098 URL https://www.sciencedirect.com/science/article/pii/0038109894006628
- [13] Hardy W N and Whitehead L A 1981 Review of Scientific Instruments 52 213–216 (Preprint eprint https://doi.org/10.1063/1.1136574) URL https://doi.org/10.1063/1.1136574
- [14] Froncisz W and Hyde J S 1982 Journal of Magnetic Resonance (1969) 47 515–521 ISSN 0022-2364 URL https://www.sciencedirect.com/science/article/pii/0022236482902219
- [15] Mourigal M, Enderle M, Kloepperpieper A, Caux J S, Stunault A and Ronnow H M 2013 Nature Physics 9 435–441 ISSN 1745-2473 URL https://doi.org/10.1038/Nphys2652
- [16] Bai X, Paddison J, Kapit E, Koohpayeh S, Wen J J, Dutton S, Savici A, Kolesnikov A, Granroth G, Broholm C, Chalker J and Mourigal M 2019 Physical Review Letters 122 097201 URL https://doi.org/10.1103/PhysRevLett.122.097201
- [17] Eisenach E R, Barry J F, Pham L M, Rojas R G, Englund D R and Braje D A 2018 Review of Scientific Instruments 89 ISSN 0034-6748 094705 (Preprint eprint https://pubs.aip.org/aip/rsi/article-pdf/doi/10.1063/1.5037465/16140529/094705_1_online.pdf) URL https://doi.org/10.1063/1.5037465
- [18] Rinard G, Quine R W, Eaton S S and Eaton G R 1993 Journal of Magnetic Resonance, Series A 105 137–144 URL https://api.semanticscholar.org/CorpusID:98346578
- [19] Koohpayeh S, Wen J J, Mourigal M, Dutton S, Cava R, Broholm C and McQueen T 2013 Journal of Crystal Growth 384 39–43 ISSN 0022-0248 URL https://www.sciencedirect.com/science/article/pii/S0022024813005915
- [20] Libersky M M, Silevitch D M and Kouki A 2019 Design of a loop-gap resonator with bimodal uniform fields using finite element analysis 2019 22nd International Conference on the Computation of Electromagnetic Fields (COMPUMAG) pp 1–4
- [21] Blanchard C, Deville A, Boukenter A, Champagnon B and Duval E 1986 Journal de Physique 47 1931–1937 ISSN 0021-1038 URL https://doi.org/10.1051/jphys:0198600470110193100
- [22] Yang Y, Babkevich P, Gaal R, Zivkovic I and Ronnow H M 2023 Cavity-magnon-polariton spectroscopy of strongly hybridized electro-nuclear spin excitations in lihof4 (Preprint eprint 2309.05051)
- [23] Rinard G A and Eaton G R 2005 Loop-gap resonators Biomedical EPR, Part B: Methodology, Instrumentation, and Dynamics Biological Magnetic Resonance ed Eaton S R, Eaton G R and Berliner L J (Springer US) pp 19–52
- [24] Poole C P 1996 Electron Spin Resonance: A Comprehensive Treatise on Experimental Techniques (Courier Corporation) ISBN 978-0-486-69444-3
- [25] Collett C A, Allão Cassaro R A and Friedman J R 2016 Phys. Rev. B 94(22) 220402 URL https://link.aps.org/doi/10.1103/PhysRevB.94.220402
- [26] Petersan P J and Anlage S M 1998 Journal of Applied Physics 84 3392–3402 ISSN 0021-8979 URL https://doi.org/10.1063/1.368498
- [27] McKenzie R D, Libersky M, Silevitch D M and Rosenbaum T F 2022 Phys. Rev. A 106(4) 043716
- [28] Yu, C. X., Zihlmann, S., Abadillo-Uriel, J. C., et al. 2023 Nat. Nanotechnol. 18, 741–746. https://doi.org/10.1038/s41565-023-01332-3